OpenStudy (anonymous):

multiply and simply rational expressions please help x^2/x^2+3 * x+5/x

OpenStudy (anonymous):

\[\frac{ x^{2} }{ x^{2} + 3 } \times \frac{ x + 5 }{ x}\]

OpenStudy (anonymous):

do u mean that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you take me step by step, I have9 questions for rational and i get some of them but some I dont

OpenStudy (anonymous):

\[\frac{ x^{2} (x+5)}{ (x^{2} + 3)5}\]

OpenStudy (anonymous):

ok ?? do you still follow me??

OpenStudy (anonymous):

yes I believe so you

OpenStudy (anonymous):

please try to the next.., and i'll see

OpenStudy (anonymous):

@tiawanad what did u get it?

OpenStudy (anonymous):

gerryliyana are you still here

OpenStudy (anonymous):

yes i'm here